Gita is an auto mechanic and runs a small repair shop. She hires one mechanic at $20,000 per year, pays a monthly rent of $5,000 towards the lease of the premises where her repair shop is located, and spends $20,000 per year on materials needed for repairing cars and trucks. She has invested $40,000 of her own savings in heavy equipment (air and strut compressors, brake lathes, heavy-duty lifts etc.) that could earn her $4,000 per year if invested elsewhere. She has been offered $40,000 per year by a competitor to work as a mechanic for them and she estimates the value of income she could earn in her spare time to be $5,000. The repair shop’s total annual revenue is $250,000. What is Gita’s yearly fixed cost in dollars? what is Gita’s yearly variable cost in dollars? what is Gita’s yearly explicit cost? what is Gita’s yearly implicit cost? what is Gita’s accounting profit? and what is Gita's yearly economic profit? and finally is Gita making the best use of her resources (time and money) in terms of economic profit?
Answer #1

Fixed cost donot depend upon the output level.

Fixed Cost=Mechanic salary+rent=20000+5000=$25000

Variable cost depend upon the voulme of business i.e. number of services.

Annual Variable Cost=Material cost=$20000

Explicit Costs include accounting costs. So,

Explicit Cost=Fixed Cost+Total variable Cost=25000+20000=$45000

Implicit Costs include the opportunity costs. So,

Implicit Cost=Salary that could been earned+interest loss+Value of spare time

Implicit Cost=40000+4000+5000=$49000

Accounting Profit=Total Revenue-Explicit Cost=250000-45000=$205,000

Economic Profit=Total Revenue-Explicit Cost-Implicit Cost

Economic Profit=250000-45000-49000=$156000

Since economic profit is highet than zero, we can say that Gita is making best use of her resources.